OpenStudy (tvinsant):

What is the solution of the following? -3y - 2y = -12 9x + 6y = -9 a. (2,1) b. no solutions c. (-2,-1) d. infinitely many solutions

OpenStudy (shinalcantara):

have you tried solving it? you can use elimination and/or substitution method to find x and y

OpenStudy (tvinsant):

i think its b. but i just need to check

OpenStudy (shinalcantara):

we'll check -3x - 2y = -12 -----equation 1 9x + 6y = -9 ----equation 2 dividing the whole of equation 2 with 3 3x + 2y = -9 ------equation 2' multiplying equation 1 with -1 3x + 2y =-12 ------equation 1' equating both equations 1' and 2' 3x + 2y + 9 = 3x + 2y + 12 3x and 2y will cancel 9 is not equal to 12 therefore there is no solution

OpenStudy (tvinsant):

Thank you for the walkthrough!